# MixCache: Mixture-of-Cache for Video Diffusion Transformer Acceleration

### v1

@@ 4.4 Adaptability
4.5 Scaling to More GPUs and Higher Resolution
@@
MixCache can be compatible with the current mainstream DiT parallel methods with a minor cache selection synchronization overhead. We integrate Ulysses parallel (Jacobs et al. 2023) to MixCache and present the latency on Wan 14B. As in Figure 10, the parallel version of MixCache still demonstrates a consistent strong scaling with increasing GPU configurations. Besides, this performance advantage is maintained across varying resolutions, demonstrating its effectiveness on high-resolution video generation tasks.
5 Conclusion
@@
To address the high-latency challenge of video DiT inference caused by its multi-step iterative denoising process, we propose MixCache, a training-free caching-based framework for efficient video DiT inference. By leveraging the redundancy in the diffusion process of different granularities and adaptively combines three-level caching (step/cfg/block), MixCache achieves comparable generation quality while significantly improving inference efficiency, outperforming existing baselines with a maximum speedup of 1.94 on Wan 14B and 1.97 on HunyuanVideo. This study establishes hybrid caching as a novel and effective approach for accelerating video DiT inference while maintaining quality.
Appendix A Appendix
A.1 Pseudo code of the MixCache framework
The execution flow of MixCache is formalized in Algorithm 1. The methodology commences with an offline profiling phase to determine the hyper-parameters (line 1). During runtime, one generation is partitioned into two distinct phases: the warm up phase and the cache enabled phase. In the warm up phase, the system exclusively executes full computations to maintain video quality. The cache enabled phase implements a dynamic switching mechanism that alternates between cached computation at three granularities and full computation, utilizing a predefined as the transition criterion (line 5). Once in the cache enabled phase, two core control mechanisms are activated: (1) the scaling strategy to regulate the frequency of cache interval, and (2) the adaptive hybrid cache decision strategy (line 23-24) that dynamically determines the optimal caching granularity for each timestep.
A.2 Further details of baseline methods
For TeaCache, we use the open-source implementation and adjust the L1_distance_thresh parameter to balance quality and speed. Teacache0.1 indicates that the value of L1_distance_thresh is set to 0.1. Increasing the L1_distance_thresh enhances acceleration efficiency. However, this improvement comes at the cost of reduced generation quality.
For BlockDance, following the official setting, we set the cache block index as 20 and set the first 40% of denoising steps as warm up steps that disable caching, and evenly divide the remaining 60% of denoising steps into several groups, each comprising steps. BlockDance4 denotes that the value equals to 4.
As for PAB, we adopt the implementation from HuggingFace Diffusers, tuning both block_skip_range and timestep_skip_range to manage the quality-speed trade-off. PAB denotes a configuration where the block_skip_range is set to 8 and the timestep_skip_range is set to .
A.3 Further details of the offline profiling process
In the offline profiling phase, we execute 100 prompts to inference to collect operational data for determining runtime hyper-parameters (, , ). As this process occurs offline, it requires only a single execution, enabling deployment of these hyper-parameters during runtime without introducing additional computational overhead. As a result, the profiled hyper-parameters , and for Wan and HunyuanVideo are 0.1, 0.05, 0.1, respectively, and those for CogVideoX are 0.1, 0.3, 0.4 respectively.
A.4 More visualization results
To comprehensively verify the method we proposed, we present additional visual presentation for each DiT model from extended VBench prompts, as indicated in Figure 11 and Figure 12. Our method maintains high-quality with a high degree of consistency and fine details with the video generated by the original models, while achieving significant acceleration.
In addition, we demonstrate the generated videos of three models using the same prompts under different MixCache configurations, including accuracy-prior and efficiency-prior configurations. As shown in Figure 13, these two configurations of MixCache show a highly consistent generation quality with the original video across different prompts and base models, reflecting its effectiveness.
A.5 A.5 Full VBench score for each dimension
Tables 3 and Tables 4 present the detailed VBench scores of Wan, HunyuanVideo and CogVideoX across various dimensions. MixCache can maintain a high VBench score, including the total score and the scores of all dimensions, indicating its advantages over baselines in terms of generation quality.
| Score | original | Teacache0.1 | Teacache0.14 | Fastercache | BlockDance4 | PAB | MixCache | MixCache |
| Wan 14B 480p | ||||||||
| total score | 84.05 | 84.01 | 83.95 | 83.40 | 83.48 | 83.00 | 83.97 | 83.90 |
| quality score | 85.04 | 84.96 | 84.85 | 84.26 | 84.29 | 83.77 | 84.89 | 84.84 |
| semantic score | 80.09 | 80.24 | 80.34 | 79.94 | 80.25 | 79.95 | 80.27 | 80.16 |
| subject consistency | 93.91 | 93.85 | 93.84 | 92.87 | 93.31 | 93.58 | 93.97 | 93.94 |
| background consistency | 97.06 | 97.08 | 97.08 | 95.11 | 95.87 | 94.46 | 97.10 | 97.17 |
| temporal flickering | 96.84 | 96.87 | 96.82 | 96.65 | 96.52 | 97.11 | 96.82 | 96.79 |
| motion smoothness | 93.76 | 93.72 | 93.86 | 94.24 | 93.76 | 94.37 | 93.89 | 93.89 |
| dynamic degree | 37.50 | 37.22 | 36.95 | 36.80 | 36.25 | 35.41 | 36.80 | 36.66 |
| aesthetic quality | 66.90 | 66.77 | 66.75 | 65.78 | 65.84 | 62.93 | 66.82 | 66.77 |
| imaging quality | 66.76 | 66.71 | 66.22 | 66.24 | 66.34 | 66.62 | 66.37 | 66.21 |
| object class | 93.09 | 93.47 | 93.01 | 92.29 | 93.64 | 92.55 | 93.20 | 93.02 |
| multiple objects | 81.16 | 79.10 | 80.79 | 81.14 | 80.66 | 80.75 | 81.10 | 80.66 |
| human action | 98.00 | 97.60 | 97.40 | 97.80 | 98.00 | 97.80 | 97.80 | 97.60 |
| color | 84.60 | 85.08 | 83.63 | 83.79 | 83.83 | 83.98 | 84.72 | 84.94 |
| spatial relationship | 78.65 | 80.53 | 80.32 | 79.84 | 81.22 | 77.41 | 79.06 | 79.08 |
| scene | 66.58 | 67.92 | 69.81 | 67.20 | 66.69 | 69.20 | 68.30 | 68.07 |
| appearance style | 77.20 | 77.02 | 76.81 | 76.56 | 76.88 | 76.77 | 76.91 | 76.88 |
| temporal style | 68.35 | 68.30 | 68.21 | 67.77 | 68.30 | 68.19 | 68.30 | 68.10 |
| overall consistency | 73.16 | 73.13 | 73.05 | 73.02 | 73.08 | 72.88 | 73.08 | 73.10 |
| Wan 14B 720p | ||||||||
| total score | 83.66 | 83.75 | 83.78 | 83.12 | 83.43 | 83.32 | 83.74 | 83.70 |
| quality score | 84.66 | 84.74 | 84.71 | 84.08 | 84.35 | 84.13 | 84.76 | 84.67 |
| semantic score | 79.69 | 79.78 | 80.05 | 79.30 | 79.77 | 80.06 | 79.68 | 79.83 |
| subject consistency | 92.45 | 92.34 | 92.41 | 91.39 | 91.95 | 92.21 | 92.34 | 92.47 |
| background consistency | 96.64 | 96.63 | 96.59 | 94.95 | 95.68 | 95.13 | 96.72 | 96.70 |
| temporal flickering | 97.17 | 97.14 | 97.11 | 97.01 | 96.84 | 97.41 | 97.14 | 97.14 |
| motion smoothness | 92.66 | 92.80 | 93.00 | 93.52 | 93.07 | 93.45 | 93.07 | 93.17 |
| dynamic degree | 37.22 | 38.05 | 38.05 | 37.09 | 37.64 | 36.80 | 37.91 | 37.36 |
| aesthetic quality | 65.67 | 65.58 | 65.48 | 64.75 | 65.01 | 63.50 | 65.76 | 65.63 |
| imaging quality | 68.46 | 68.26 | 67.99 | 67.82 | 68.06 | 68.34 | 68.00 | 67.90 |
| object class | 92.01 | 91.71 | 91.65 | 90.76 | 91.61 | 91.52 | 92.14 | 92.36 |
| multiple objects | 78.19 | 78.93 | 78.26 | 76.89 | 79.13 | 77.94 | 78.99 | 78.75 |
| human action | 98.20 | 98.20 | 98.00 | 97.80 | 98.20 | 98.20 | 98.00 | 97.80 |
| color | 86.20 | 86.69 | 88.57 | 85.79 | 87.59 | 86.14 | 86.58 | 85.36 |
| spatial relationship | 76.09 | 75.93 | 76.32 | 75.85 | 75.68 | 78.82 | 75.01 | 76.58 |
| scene | 68.60 | 69.03 | 70.04 | 69.82 | 68.27 | 71.31 | 68.73 | 70.02 |
| appearance style | 76.07 | 75.72 | 75.79 | 75.26 | 75.65 | 75.23 | 75.79 | 75.72 |
| temporal style | 67.77 | 67.97 | 67.66 | 67.53 | 67.86 | 67.72 | 67.86 | 67.86 |
| overall consistency | 74.12 | 73.85 | 74.18 | 73.96 | 73.98 | 73.65 | 73.98 | 74.01 |
| Score | original | Teacache0.1 | BlockDance4 | PAB | MixCache | MixCache |
| HunyuanVideo 540p | ||||||
| total score | 81.13 | 80.87 | 80.93 | 80.64 | 81.05 | 80.98 |
| quality score | 82.66 | 82.32 | 82.28 | 82.21 | 82.56 | 82.51 |
| semantic score | 75.04 | 75.05 | 75.52 | 74.36 | 75.00 | 74.85 |
| subject consistency | 92.86 | 92.75 | 92.63 | 92.77 | 92.90 | 92.91 |
| background consistency | 95.37 | 95.36 | 94.75 | 95.11 | 95.46 | 95.42 |
| temporal flickering | 98.22 | 98.25 | 97.65 | 98.11 | 98.22 | 98.25 |
| motion smoothness | 96.47 | 96.36 | 96.12 | 96.47 | 96.54 | 96.60 |
| dynamic degree | 31.25 | 30.28 | 30.28 | 30.14 | 30.84 | 30.70 |
| aesthetic quality | 61.41 | 61.46 | 60.96 | 60.63 | 61.41 | 61.41 |
| imaging quality | 61.71 | 60.65 | 62.41 | 61.11 | 61.26 | 61.05 |
| object class | 86.11 | 87.50 | 85.97 | 85.55 | 86.12 | 86.31 |
| multiple objects | 68.16 | 67.41 | 68.43 | 67.03 | 66.94 | 66.28 |
| human action | 97.20 | 97.20 | 97.20 | 97.20 | 97.20 | 97.60 |
| color | 85.09 | 85.81 | 86.13 | 86.56 | 86.71 | 85.69 |
| spatial relationship | 61.29 | 61.15 | 66.04 | 59.57 | 61.45 | 61.69 |
| scene | 60.44 | 59.72 | 58.74 | 57.82 | 59.83 | 59.64 |
| appearance style | 77.72 | 77.72 | 77.86 | 77.51 | 77.48 | 77.37 |
| temporal style | 66.46 | 66.87 | 66.37 | 65.71 | 66.48 | 66.43 |
| overall consistency | 72.86 | 72.09 | 72.94 | 72.28 | 72.77 | 72.66 |
| CogVideoX 5B 480p | ||||||
| total score | 80.89 | 80.15 | 74.34 | 78.67 | 80.10 | 80.15 |
| quality score | 82.20 | 81.22 | 76.12 | 80.35 | 81.42 | 81.44 |
| semantic score | 75.61 | 75.87 | 67.24 | 71.95 | 74.80 | 74.99 |
| subject consistency | 92.55 | 91.92 | 89.20 | 91.75 | 92.00 | 92.04 |
| background consistency | 94.48 | 94.07 | 93.45 | 94.08 | 94.27 | 94.16 |
| temporal flickering | 93.90 | 93.66 | 91.96 | 95.28 | 93.82 | 93.80 |
| motion smoothness | 92.80 | 91.90 | 88.71 | 93.83 | 92.18 | 92.04 |
| dynamic degree | 38.47 | 35.84 | 15.14 | 33.75 | 36.25 | 36.53 |
| aesthetic quality | 60.69 | 59.95 | 56.00 | 58.37 | 60.02 | 59.99 |
| imaging quality | 61.44 | 60.60 | 60.31 | 55.24 | 60.71 | 60.80 |
| object class | 87.41 | 86.16 | 74.37 | 81.63 | 86.11 | 86.28 |
| multiple objects | 65.09 | 66.01 | 44.41 | 56.25 | 64.27 | 64.77 |
| human action | 98.20 | 97.80 | 96.40 | 96.80 | 98.00 | 98.20 |
| color | 86.84 | 87.16 | 76.28 | 88.84 | 87.30 | 87.66 |
| spatial relationship | 55.03 | 57.60 | 49.66 | 48.21 | 54.31 | 54.23 |
| scene | 66.46 | 65.60 | 50.84 | 59.38 | 62.79 | 63.16 |
| appearance style | 81.13 | 81.66 | 81.62 | 80.18 | 81.17 | 81.10 |
| temporal style | 66.87 | 66.62 | 61.59 | 64.31 | 65.82 | 65.99 |
| overall consistency | 73.46 | 74.18 | 70.00 | 71.92 | 73.46 | 73.52 |

### v2

@@ 4.4. Adaptability
4.5. Latency Breakdown
Figure 10 shows the latency breakdown of MixCache on three models. Among them, “Full” indicates the time for performing full computations at adaptive intervals after entering the cache enabled phase. “Overhead” is primarily used for determining cache intervals, calculating values, and selecting cache modes for each step. As step-level caching skips all the computation of that step, it is not included in the breakdown. As can be seen, MixCache’s performance benefit primarily comes from significantly reducing full computation time with minimal overhead.
4.6. Hyperparameter Selection
We conduct experiments under different hyperparameter configurations on a wan 14B 480p with 200 prompts. Figure 11 plots the generated quality (indicated by PSNR ↑) and latency for: (a) varying with (, )=(0.05,0.1) and (b) varying (, ) with =0.1. As can be seen, under different configurations, there is a trade-off between quality and latency. Specifically, Figure 11(a) shows that, when fixed (, )=(0.05,0.1), while =0.08 or 0.05 has a slight improvement in generated quality compared to =0.1, it suffers from a significant disadvantage in latency. This phenomenon can also be seen in Figure 11(b). Therefore, considering the trade-off between quality and speed, we choose the configuration with =0.1 and (, )=(0.05,0.1).
4.7. Block Candidate Selection
We experiment the effects of different block candidates on quality, latency, and memory overhead. As shown in Table 3, the more block candidates, the more extra memory overhead will be incurred, as more block outputs need be saved for caching. However, a larger number of block candidates does not necessarily lead to improvements in generation quality and speed. In addition, if there are too few block candidates, the generation quality will drop significantly, since the block level caching cannot capture the different features of different blocks. Therefore, we select (10, 20, 30) as the default block configuration, which has the optimal performance in terms of quality and speed in our experiments, and does not incur excessive memory overhead.
| Block Candidate | LPIPS ↓ | PSNR ↑ | SSIM ↑ | Latency | Mem. |
| (20) | 0.093 | 25.04 | 0.803 | 483 s | 0.64 GB |
| (10,30) | 0.089 | 25.22 | 0.821 | 470 s | 1.27 GB |
| (10,20,30) | 0.079 | 25.91 | 0.858 | 465 s | 1.89 GB |
| (10,15,20,25,30,35) | 0.080 | 25.86 | 0.849 | 468 s | 3.77 GB |
4.8. Scaling to Multi-GPUs and Higher Resolution
@@
MixCache can be compatible with the current mainstream DiT parallel methods with a minor cache selection synchronization overhead. We integrate Ulysses parallel (Jacobs et al., 2023) to MixCache and present the latency on Wan 14B. As in Figure 12, the parallel version of MixCache still demonstrates a consistent strong scaling with increasing GPU configurations. Besides, this performance advantage is maintained across varying resolutions, demonstrating its effectiveness on high-resolution video generation tasks.
5. Conclusion
@@
To address the high-latency challenge of video DiT inference caused by its multi-step iterative denoising process, we propose MixCache, a training-free caching-based framework for efficient video DiT inference. By leveraging the redundancy in the diffusion process of different granularities and adaptively combines three-level caching (step/cfg/block), MixCache achieves comparable generation quality while significantly improving inference efficiency, outperforming existing baselines with a maximum speedup of 1.94 on Wan 14B and 1.97 on HunyuanVideo. Our work establishes hybrid caching as a novel and effective approach for accelerating video DiT inference, supporting responsive multimedia content generation.